OpenStudy (anonymous):

Find the Taylor series with a = 0 of sin(x) Find the Taylor series with a = 0 of sin(x) @Mathematics

OpenStudy (anonymous):

x-(x^3/3!)+((x^5/5!)-(x^7)/7!+.......

OpenStudy (anonymous):

express it as an infinite sum :-D

OpenStudy (anonymous):

or an infinite series*

OpenStudy (anonymous):

ok i got ur point

OpenStudy (anonymous):

u mean in terms of summation

OpenStudy (anonymous):

yes something like \[\sum_{n=0}^{\infty}\]

OpenStudy (anonymous):

yes i am writing

OpenStudy (anonymous):

Wow this crap is way over my head...

OpenStudy (anonymous):

\[\sum_{0}^{\infty\[\frac{x ^{2n+1}}{(2n+1)!}\]}\]

OpenStudy (anonymous):

\[\sum_{0}^{\infty}\]\[\frac{x^2n+1}{(2n+1)!}\]

OpenStudy (anonymous):

this is right and as required

OpenStudy (anonymous):

\[\sum_{n=0}^{\infty}{(-1)^n\frac{x^{2n+1}}{(2n+1)!}}\]

OpenStudy (anonymous):

yes i forgot to put -1 for alternating sign